OpenStudy (anonymous):

Which of the following equation is written in point -slope form a. y=3x + 4 b. 2x+5y=5 C. y-5=2(x-1

OpenStudy (anonymous):

a. y=3x+4 Point slope form is like this Exp: y=mx+c

OpenStudy (osanseviero):

Point slope is: y - y1 = m (x-x1) It is named point slope because you have the coordinates of a point P (x1,y1) and a slope (m)

OpenStudy (osanseviero):

So from that, check the three

OpenStudy (anonymous):

OH MY GOD SRRY xDD

OpenStudy (osanseviero):

You can check this link if you have any doubts: http://www.mathsisfun.com/algebra/line-equation-point-slope.html